OpenStudy (anonymous):

What is the answer for 150r divided by sqrt(r^3)

OpenStudy (anonymous):

\[\frac{150r}{\sqrt{r^3}}=\frac{150r}{r\sqrt{r}}\] is a start

OpenStudy (anonymous):

then cancel an \(r\)

OpenStudy (anonymous):

Which r though?

OpenStudy (anonymous):

\[\frac{ 150 }{ \sqrt{r} }\]

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

if you are a stickler for writing without a radical in the denominator, you should write \[\frac{150\sqrt{r}}{r}\] or leave it alone